OpenStudy (anonymous):

Choose the point which is a solution to the system of inequalities: 4x + 3y > 12 x – y –2 Choices: A.(2, 1) B.(–5, –6) C.(–4, 9) D.(3, 7) Tell me how to solve it and the answer please

OpenStudy (anonymous):

what's the second inequality?

OpenStudy (anonymous):

the points are (x,y). So for ur first point, 2 is x and 1 is y. plug them into your first inequality and see if they work.

OpenStudy (anonymous):

oh it's less than or equal to

myininaya (myininaya):

\[x-y \le -2 ?\]

OpenStudy (anonymous):

then you should plug it into your second and see if it works there. so it's x-y is less than or equal to 2?

myininaya (myininaya):

\[x-y \le 2?\]

OpenStudy (anonymous):

yeah thats it

myininaya (myininaya):

the second one?

OpenStudy (anonymous):

yeah but instead of the 2 its -2

myininaya (myininaya):

so the first one?

myininaya (myininaya):

\[x-y \le -2\]

OpenStudy (anonymous):

k. so ,first equation w/first point- 4(2)+3(1)>12 8+3>12 11>12 no so the first point isn't it

myininaya (myininaya):

go through the choices and see what which of the choices satisfy the system of inequalities

OpenStudy (anonymous):

second point- 4(-5)+3(-6)>12 -20-18>12 -38>12 no again

OpenStudy (anonymous):

third- 4(-4)+3(9)>12 -16+27>12 11>12 no

myininaya (myininaya):

he is saying no this point won't work because 11>12 is not true and -38>12 is not true so he has pointed out so far (2,1) and (-5,-6) will not work

OpenStudy (anonymous):

it has to be the last one by process of elimination, but we'll work it through. 4(3)+3(7)>12 12+21>12 33>12 yes

OpenStudy (anonymous):

hello myininaya

myininaya (myininaya):

hey

OpenStudy (anonymous):

ok i got it thank you :)

OpenStudy (anonymous):

i wanted to tell you that yesterday some one wanted to solve a quadric equation

myininaya (myininaya):

needhelp you have to see if it works for both inequalities

OpenStudy (anonymous):

and they say "can we use the Method of Myininaya" that i heard about?"

myininaya (myininaya):

lol

OpenStudy (anonymous):

method of myininaya! really

OpenStudy (anonymous):

i got another question

OpenStudy (anonymous):

don't forget to put it through the second system too- x-y<or equal to 2 3-7<equal2 -4<equal 2 yup

OpenStudy (anonymous):

sure! we'll help!

OpenStudy (anonymous):

ok here it is

myininaya (myininaya):

so it still would for -4<equal -2 which ever one you had

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Choose the system of inequalities that matches the graph below

OpenStudy (anonymous):

and here are the choices

OpenStudy (anonymous):

A)2x – 3y > –27 x + 5y < 15 B)2x – 3y < –27 x + 5y < 15 C)2x – 3y –27 x + 5y 15 D)2x – 3y –27 x + 5y 15

OpenStudy (anonymous):

inequality signs are missing last 2 answers

OpenStudy (anonymous):

pretty sure one of the lines is y>2/3 x+9... now just need to see what form it's in here...

OpenStudy (anonymous):

I mean y<

OpenStudy (anonymous):

oh choice 2 is 2x-3 greater than or equal to -27 and x+5y less than or equal to 15

myininaya (myininaya):

its look at the lines first lets look at the one with positive slope it has y intercept 9 and rise over run is 2/3 so that line would be y=2x/3+9 now lets look at the second line that has negative slope it has y-intercept 3 and rise over run is -1/5 so that line would be y=-x/5+3 now lets look at the inequality part

OpenStudy (anonymous):

y<-1/5 x+3 is ze second...

OpenStudy (anonymous):

choice 3 is 2x – 3y Less than or equal to –27 and x + 5y less than or equal to 15

myininaya (myininaya):

both lines are broken so we don't have = signs both should have less than sign and gandalf gots it

myininaya (myininaya):

now put in the form they have it in and you will find your answers

OpenStudy (anonymous):

yeah i just gave you the signs for each

myininaya (myininaya):

oops that first one should have a less than sign

myininaya (myininaya):

-2x+3y<27

myininaya (myininaya):

y<2x/3+9 3y<2x+27 -2x+3y<27

OpenStudy (anonymous):

so the answer is B?

myininaya (myininaya):

2x-3y>-27 x+5y<15

OpenStudy (anonymous):

oooooh its A

OpenStudy (anonymous):

thank you so much again :)